# Locus: Agentic Predicate Synthesis for Directed Fuzzing

### v2

@@ 3.4. Validation
The validation step is critical to preserving the correctness of the instrumented program and ensuring that the inserted predicate $\phi_{l}$ maintains the fuzzing admissibility. As shown in Algorithm 1 lines 7 to 9, the validator takes as input the candidate predicate $\phi$, the vulnerability canary $\psi$, and the program $P$, and validates whether the predicate is both syntactically and semantically correct. If the predicate passes both checks, Locus will refine the predicate by exploring potentially better program points closer to the program entries (Section 3.3). If it fails, Locus will self-reflect and attempt to regenerate the predicate, using diagnostic feedback collected from the validator.
@@
Semantic validation  The second component is semantic validation, which confirms that the predicate $\phi$ strictly relaxes $\psi$. It ensures that the predicate does not reject any execution paths that could reach the vulnerability (Theorem 4). As we demonstrated in Definition 3, such validation requires enumerating all possible program inputs, which cannot be done within an acceptable time budget. Therefore, we utilize symbolic execution to find counterexamples of the relaxation.
@@
It is natural to employ symbolic execution as a formal checker for Theorem 5 to ensure fuzzing admissibility. However, it often incurs substantial overhead by exploring irrelevant execution paths(Cadar et al., 2008), e.g., branches that are unrelated to either the synthesized predicate or the target canary. This excessive path exploration can lead to prohibitively long validation time, incurring additional overhead of deploying Locus. To mitigate this inefficiency, we adopt a strategy inspired by Chopper(Trabish et al., 2018) that skips specified parts of the code and targets only the exploration of paths according to our selection. Specifically, we perform a lightweight reachability analysis on the CFG and prune nodes that are not reachable from either the predicate or the canary. Locus then initiates symbolic execution to explore the path between the $\neg\phi$ and $\psi$.
@@ 4. Evaluation
Table 2. TTE for each vulnerability in the Magma benchmark across different directed fuzzers. T.O. indicates that the fuzzer cannot find the vulnerability within 24 hours. $\varnothing$ indicates that the fuzzer either could not compile the target program or the preprocessing step exceeded 24 hours.
Vul ID AFLGo SelectFuzz Beacon Titan
origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$
PNG003 4 4 1.0 1 3 3 1.0 1 5 5 1.0 1 8 8 1.0 1
PNG006 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
PNG007 72536 9052 8.0 ¡0.01 58770 8537 6.9 0.04 7764 1659 4.7 ¡0.01 17561 11954 1.5 0.04
SND001 T.O. 419 206.5 ¡0.01 7764 5 1725.4 ¡0.01 432 13 33.2 ¡0.01 115 8 14.4 ¡0.01
SND005 1015 205 5.0 ¡0.01 102 4 23.2 ¡0.01 95 16 5.9 ¡0.01 30412 1046 29.1 ¡0.01
SND006 T.O. 3899 22.2 ¡0.01 8519 12 709.9 ¡0.01 T.O. T.O. - - 26171 905 28.9 ¡0.01
SND007 T.O. 3468 24.9 ¡0.01 7785 62 126.4 ¡0.01 T.O. T.O. - - 310 28 11.1 ¡0.01
SND017 4453 1021 4.4 ¡0.01 3235 1518 2.1 ¡0.01 4861 58 83.8 ¡0.01 40955 256 160.0 ¡0.01
SND020 3805 1291 2.9 ¡0.01 694 542 1.3 ¡0.01 5683 41 138.6 ¡0.01 70432 604 116.7 ¡0.01
SND024 T.O. 2577 33.5 ¡0.01 5957 18 332.8 0.05 T.O. T.O. - - 1074 8 127.8 ¡0.01
TIF002 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF005 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF006 80998 17774 4.6 ¡0.01 T.O. T.O. - - 35493 6715 5.3 0.02 66692 65412 1.0 0.04
TIF007 18513 2107 8.8 ¡0.01 1410 413 3.4 ¡0.01 220 76 2.9 ¡0.01 146 109 1.3 ¡0.01
TIF009 32556 16767 1.9 0.04 12067 4144 2.9 ¡0.01 T.O. T.O. - - T.O. T.O. - -
TIF012 47424 18689 2.5 0.02 9118 4787 1.9 ¡0.01 18782 12274 1.5 0.04 2922 2620 1.1 ¡0.01
TIF014 66000 8134 8.1 ¡0.01 58851 7833 7.5 ¡0.01 19035 9672 2.0 0.02 44384 14360 3.1 ¡0.01
LUA004 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - 69594 67442 1.0 ¡0.01
XML001 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
XML003 T.O. T.O. - - T.O. T.O. - - 44056 13295 3.3 ¡0.01 53689 36365 1.5 0.32
XML009 T.O. 77513 1.1 0.6 T.O. 56 1540.1 ¡0.01 32585 6748 4.8 ¡0.01 17942 13879 1.3 0.03
XML017 17 8 2.1 ¡0.01 7 9 0.7 ¡0.01 10 15 0.7 0 23 13 1.8 ¡0.01
SSL001 T.O. T.O. - - 72078 62637 1.2 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
SSL003 222 231 1.0 0.04 73 69 1.1 0.02 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP004 165 269 0.6 ¡0.01 11 11 1.0 1 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP009 99 203 0.5 0.04 185 153 1.2 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
PHP011 $\varnothing$ $\varnothing$ - - 26 23 1.1 ¡0.01 $\varnothing$ $\varnothing$ - - $\varnothing$ $\varnothing$ - -
SQL018 49520 30855 1.6 0.03 53397 6748 7.9 ¡0.01 T.O. T.O. - - 78238 42275 1.9 0.04
[…]
Table 3. TTE for each vulnerability in the Magma benchmark across different coverage-guided fuzzers. T.O. indicates that the fuzzer cannot find the vulnerability within 24 hours. $\varnothing$ indicates that the fuzzer either could not compile the target program or the preprocessing step exceeded 24 hours.
Vul ID AFL++ AFL MOPT Fox
origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$ origin Locus ratio $p$
PNG003 12 8 1.5 0.04 3 3 1.0 1.00 4 4 1.0 1.00 5 5 1.0 1.00
PNG006 106 63 1.7 0.16 T.O. T.O. - - T.O. T.O. - - 6978 1582 4.4 ¡0.01
PNG007 53104 41101 1.3 0.05 55351 18358 3.0 ¡0.01 42811 37522 1.1 0.07 4651 4009 1.2 0.05
SND001 451 19 24.4 ¡0.01 75940 338 224.4 ¡0.01 285 12 23.6 ¡0.01 447 27 16.6 ¡0.01
SND005 1233 193 6.4 ¡0.01 796 9 88.4 ¡0.01 53 9 6.3 ¡0.01 82 15 5.5 ¡0.01
SND006 7026 24 290.3 ¡0.01 T.O. 2251 38.4 ¡0.01 328 17 19.6 ¡0.01 15422 717 21.5 ¡0.01
SND007 810 34 23.6 ¡0.01 T.O. 3163 27.3 ¡0.01 360 16 22.4 ¡0.01 3226 129 25.0 ¡0.01
SND017 841 196 4.3 ¡0.01 3162 833 3.8 ¡0.01 33 20 1.7 ¡0.01 2117 178 11.9 ¡0.01
SND020 640 204 3.1 0.08 2916 1266 2.3 ¡0.01 326 103 3.2 ¡0.01 2068 176 11.8 ¡0.01
SND024 361 18 19.6 ¡0.01 T.O. 2391 36.1 ¡0.01 326 14 23.4 ¡0.01 2642 32 82.6 ¡0.01
TIF002 72896 69963 1.0 0.85 T.O. T.O. - - T.O. T.O. - - T.O. T.O. - -
TIF005 420 260 1.6 0.04 T.O. T.O. - - T.O. T.O. - - 4744 990 4.8 ¡0.01
TIF006 970 336 2.9 ¡0.01 52926 25366 2.1 0.02 39810 45124 0.9 0.44 22380 7607 2.9 ¡0.01
TIF007 50 26 2.0 0.04 15022 1172 12.8 0.04 155 56 2.8 ¡0.01 8868 290 30.6 ¡0.01
TIF009 17483 5234 3.3 ¡0.01 19302 14098 1.4 0.01 26049 15622 1.7 0.05 7696 3406 2.3 ¡0.01
TIF012 1731 1122 1.5 0.02 28884 16318 1.8 0.09 3639 953 3.8 0.02 1257 637 2.0 ¡0.01
TIF014 2555 682 3.7 ¡0.01 66072 7303 9.0 ¡0.01 1660 788 2.1 ¡0.01 4666 3275 1.4 ¡0.01
LUA004 20046 11939 1.7 0.03 44954 22434 2.0 0.03 12401 2098 5.9 ¡0.01 35998 6937 5.2 ¡0.01
XML001 3720 1330 2.8 ¡0.01 T.O. T.O. - - 37726 16801 2.2 0.03 T.O. T.O. - -
XML003 2373 1394 1.7 0.04 T.O. T.O. - - 56501 18636 3.0 0.03 12252 6844 1.8 0.04
XML009 2668 1452 1.8 ¡0.01 T.O. 22753 3.8 0.06 706 670 1.1 0.65 10827 8766 1.2 0.04
XML017 52 15 3.5 ¡0.01 26 7 3.7 ¡0.01 6 16 0.4 ¡0.01 77 35 2.2 ¡0.01
SSL001 23533 1583 14.9 ¡0.01 81838 T.O. 0.9 - 43952 16033 2.7 ¡0.01 66778 17764 3.8 ¡0.01
SSL003 193 178 1.1 ¡0.01 90 63 1.4 ¡0.01 96 79 1.2 ¡0.01 225 199 1.1 0.33
PHP004 66156 34974 1.9 0.03 82 89 0.9 0.03 760 185 4.1 0.03 $\varnothing$ $\varnothing$ - -
PHP009 11890 5677 2.1 0.02 143 179 0.8 0.21 650 596 1.1 0.91 $\varnothing$ $\varnothing$ - -
PHP011 1626 934 1.7 0.06 108 76 1.4 0.06 19 13 1.5 0.04 $\varnothing$ $\varnothing$ - -
SQL018 10355 4961 2.1 0.04 71038 32073 2.2 0.04 22656 28101 0.8 0.56 15498 8159 1.9 ¡0.01
@@ 4.1. Setup
Dataset  We evaluate Locus on the Magma fuzzing benchmark (Hazimeh et al., 2020), which includes a diverse set of real-world vulnerabilities selected from nine popular open-source software projects. We also consider popular libraries in the wild, e.g., libarchive, to evaluate the capabilities of Locus in finding real-world vulnerabilities (see Section 4.6).
Baselines  We select eight fuzzers (see Table 4) as our baseline, covering both the state-of-the-art directed and coverage-guided ones. For each fuzzer, we use the latest stable version available at the time of writing.
For directed fuzzers, we consider AFLGo(Böhme et al., 2017), SelectFuzz(Luo et al., 2023), Beacon(Huang et al., 2022), and Titan(Huang et al., 2024a). We do not include a comprehensive set of directed fuzzers due to the lack of publicly available implementations and the challenges of integrating them into Magma.
We also consider coverage-guided fuzzers, including the widely used AFL(Zalewski, 2020) and AFL++(Fioraldi et al., 2020). To make the evaluation more comprehensive, we include two additional fuzzers. MOPT(Lyu et al., 2019) enhances fuzzing with optimized mutation strategies, and Fox(She et al., 2024) optimizes the seed scheduling as online stochastic control.
Note that Locus alone is not a standalone fuzzer, but focuses exclusively on code transformations for the target software. Therefore, it is agnostic to the fuzzer implementations and can complement any fuzzers (Figure 2).
@@
Metrics  We follow the existing directed fuzzing approaches by adopting the Time To Exposure (TTE) to measure the performance of the baseline fuzzers and the improvement introduced by integrating Locus. TTE measures the time taken by a fuzzer to find the input that satisfies the canary condition. To mitigate the inherent randomness introduced by fuzzing, we follow (Hazimeh et al., 2020) by executing 10 independent fuzzing trials per vulnerability sample and report the average TTE. We also employ the Mann-Whitney U Test(Mann and Whitney, 1947) to demonstrate the statistical significance ($p$-value) of the results.
@@
We implement Locus using the PydanticAI framework (Colvin, 2025). The synthesizer’s tooling is primarily built on top of Multiplier (Goodman, 2025). We leverage the SVF static analysis framework (Sui and Xue, 2016) to perform a lightweight reachability analysis and use KLEE (Cadar et al., 2008) as the symbolic execution engine to perform semantic validation. We set the LLM generation temperature to zero to avoid non-determinism. We use o3-mini-2025-01-31 as the default model with reasoning level to medium, but we also evaluate other LLMs in Section 4.4.
@@ 4.2. RQ1: Vulnerabilities Reproduction
We apply Locus on both directed fuzzers and coverage-guided fuzzers, and compare the performance w/ and w/o Locus. The results are shown in Table 2 and Table 3.
@@ 4.3. RQ2: Cost Analysis
Time cost  In Section 4.2, we focused on measuring TTE, i.e., the time required to trigger the target canary during fuzzing. However, only focusing on the fuzzing phase can be misleading when evaluating the overall effectiveness of a directed fuzzer. Many directed fuzzers perform expensive static analysis on the target program before starting fuzzing (Böhme et al., 2017; Huang et al., 2024a, b). Likewise, Locus requires additional preprocessing steps, including codebase indexing, symbolic validation of predicates, and agentic predicate synthesis. While such preprocessing costs are often arguably a one-time effort that can be amortized during fuzzing, in practice, we observe that some approaches incur excessive analysis time, sometimes even longer than the time required to discover the bug by the fuzzer.
@@
LLM token cost  We assess the feasibility of deploying Locus in practice from a financial perspective, focusing on the monetary cost of LLM token usage. Table 5 shows the token cost for our Locus workflow. Locus takes 457k tokens (equivalent to \$0.72 USD) to generate predicates for one sample in the Magma benchmark on average. The monetary token costs show the potential for Locus to act as an affordable step in assisting fuzzing.
@@ 4.4. RQ3: Ablations
Table 6. TTEs by ablating different designs and models.
@@
Abalate different designs
@@
Abalate different LLMs
@@
w/ o3-mini 41101 19 1122 682 4961
w/ DR1 45104 53 992 1439 5433
w/ Gemini 38215 130 1517 2454 5274
@@
w/ O3-mini 8537 5 4787 7833 6748
w/ R1 23699 6 3562 3518 7023
w/ Gemini 7020 12 7206 40332 3921
@@
Table 6 shows that the design choices in Locus consistently improves the performance. Specifically, we observe that while predicates generated under the base setting may occasionally accelerate fuzzing, they are not consistently beneficial. Without validation, Locus sometimes generates false predicates to the program, as evidenced by the cases PNG007, TIF014, and SQL018.

### v3

@@ 3.4. Validation
The validation step is critical to preserving the correctness of the instrumented program and ensuring that the inserted predicate $\phi_{l}$ maintains the instrumented program’s fuzzing admissibility. As shown in Algorithm 1 lines 7 to 9, the validator takes as input the candidate predicate $\phi$, the vulnerability canary $\psi$, and the program $P$, and validates whether the predicate is both syntactically and semantically correct. If the predicate passes both checks, Locus will refine the predicate by exploring potentially better program points closer to the program entries (Section 3.3). If it fails, Locus will self-reflect and attempt to regenerate the predicate, using diagnostic feedback collected from the validator.
@@
Semantic validation  The second component is semantic validation, which confirms that the predicate $\phi$ strictly relaxes $\psi$. It ensures that the predicate does not reject any execution paths that can reach the vulnerability (Theorem 4). As we demonstrated in Definition 3, such validation requires enumerating all possible program inputs, which cannot be done within an acceptable time budget. Therefore, we utilize symbolic execution to find counterexamples of the relaxation.
@@
It is natural to employ symbolic execution as a formal checker for Theorem 5 to ensure fuzzing admissibility. However, it often incurs substantial overhead by exploring irrelevant execution paths(Cadar et al., 2008), e.g., branches that are unrelated to either the synthesized predicate or the target canary. This excessive path exploration can lead to prohibitively long validation time, incurring additional overhead of deploying Locus. To mitigate this inefficiency, we adopt a strategy inspired by Chopper(Trabish et al., 2018) to skip unrelated paths and target only the exploration of paths according to our selection. Specifically, we perform a lightweight reachability analysis on the CFG and prune nodes that are not reachable from either the predicate or the canary. Locus then initiates symbolic execution to explore the path between the $\neg\phi$ and $\psi$. Overall, the semantic validation offers a sound guarantee that the relaxation is harmless, but it remains unable to measure the effectiveness of the generated predicates.
@@ 4. Evaluation
Table 2. TTE for each vulnerability in the Magma benchmark across different fuzzers. T.O. indicates that the fuzzer cannot find the vulnerability within 24 hours. $\varnothing$ indicates that the fuzzer either could not compile the target program or the preprocessing step exceeded 24 hours. Numbers in bold indicate statistical significance ($p\leq 0.05$).
Vul ID AFLGo SelectFuzz Beacon Titan AFL++ AFL MOPT Fox
Origin Locus Origin Locus Origin Locus Origin Locus Origin Locus Origin Locus Origin Locus Origin Locus
PNG003 4 4 3 3 5 5 8 8 12 8 3 3 4 4 5 5
PNG006 T.O. T.O. T.O. T.O. T.O. T.O. T.O. T.O. 106 63 T.O. T.O. T.O. T.O. 6978 1582
PNG007 72536 9052 58770 8537 7764 1659 17561 11954 53104 41101 55351 18358 42811 37522 4651 4009
SND001 T.O. 419 7764 5 432 13 115 8 451 19 75940 338 285 12 447 27
SND005 1015 205 102 4 95 16 30412 1046 1233 193 796 9 53 9 82 15
SND006 T.O. 3899 8519 12 T.O. T.O. 26171 905 7026 24 T.O. 2251 328 17 15422 717
SND007 T.O. 3468 7785 62 T.O. T.O. 310 28 810 34 T.O. 3163 360 16 3226 129
SND017 4453 1021 3235 1518 4861 58 40955 256 841 196 3162 833 33 20 2117 178
SND020 3805 1291 694 542 5683 41 70432 604 640 204 2916 1266 326 103 2068 176
SND024 T.O. 2577 5957 18 T.O. T.O. 1074 8 361 18 T.O. 2391 326 14 2642 32
TIF002 T.O. T.O. T.O. T.O. T.O. T.O. T.O. T.O. 72896 69963 T.O. T.O. T.O. T.O. T.O. T.O.
TIF005 T.O. T.O. T.O. T.O. T.O. T.O. T.O. T.O. 420 260 T.O. T.O. T.O. T.O. 4744 990
TIF006 80998 17774 T.O. T.O. 35493 6715 66692 65412 970 336 52926 25366 39810 45124 22380 7607
TIF007 18513 2107 1410 413 220 76 146 109 50 26 15022 1172 155 56 8868 290
TIF009 32556 16767 12067 4144 T.O. T.O. T.O. T.O. 17483 5234 19302 14098 26049 15622 7696 3406
TIF012 47424 18689 9118 4787 18782 12274 2922 2620 1731 1122 28884 16318 3639 953 1257 637
TIF014 66000 8134 58851 7833 19035 9672 44384 14360 2555 682 66072 7303 1660 788 4666 3275
LUA004 T.O. T.O. T.O. T.O. T.O. T.O. 69594 67442 20046 11939 44954 22434 12401 2098 35998 6937
XML001 T.O. T.O. T.O. T.O. T.O. T.O. T.O. T.O. 3720 1330 T.O. T.O. 37726 16801 T.O. T.O.
XML003 T.O. T.O. T.O. T.O. 44056 13295 53689 36365 2373 1394 T.O. T.O. 56501 18636 12252 6844
XML009 T.O. 77513 T.O. 56 32585 6748 17942 13879 2668 1452 T.O. 22753 706 670 10827 8766
XML017 17 8 7 9 10 15 23 13 52 15 26 7 6 16 77 35
SSL001 T.O. T.O. 72078 62637 $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ 23533 1583 81838 T.O. 43952 16033 66778 17764
SSL003 222 231 73 69 $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ 193 178 90 63 96 79 225 199
PHP004 165 269 11 11 $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ 66156 34974 82 89 760 185 $\varnothing$ $\varnothing$
PHP009 99 203 185 153 $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ 11890 5677 143 179 650 596 $\varnothing$ $\varnothing$
PHP011 $\varnothing$ $\varnothing$ 26 23 $\varnothing$ $\varnothing$ $\varnothing$ $\varnothing$ 1626 934 108 76 19 13 $\varnothing$ $\varnothing$
SQL018 49520 30855 53397 6748 T.O. T.O. 78238 42275 10355 4961 71038 32073 22656 28101 15498 8159
@@ 4.1. Setup
Dataset  We evaluate Locus on the Magma fuzzing benchmark (Hazimeh et al., 2020), which includes a diverse set of real-world vulnerabilities selected from nine popular open-source software projects. We also consider popular libraries in the wild, e.g., VLC, to evaluate the capabilities of Locus in finding real-world vulnerabilities (see Section 4.6). Each vulnerability in Magma is represented by a canary statement (Section 2.2), placed within the target function (Figure 1). The canary condition is manually curated to specify the state necessary to trigger the vulnerability and logs the time when it is satisfied. Note that while our evaluation assumes the canary is provided, Locus can be extended to other forms of targets, including vulnerability reports, patches, or static analysis results, by introducing an additional processing step (Section 4.6).
Baselines  We select eight fuzzers (See Table 3) as our baseline, covering both the state-of-the-art directed and coverage-guided ones. For each fuzzer, we use the latest stable version available at the time of writing. Note that Locus alone is not a standalone fuzzer, but focuses exclusively on code transformations for the target software. Therefore, it is agnostic to fuzzer implementations and can complement any fuzzers (Figure 2).
@@
Metrics  We follow the existing directed fuzzing approaches by adopting the Time To Exposure (TTE) to measure the performance of the baseline fuzzers and the improvement introduced by integrating Locus. In the Magma benchmark, TTE measures the time taken by a fuzzer to find the input that satisfies the canary condition. To mitigate the inherent randomness introduced by fuzzing, we follow Hazimeh et al. (2020) by executing 10 independent fuzzing trials per vulnerability sample and report the average TTE. We also employ the Mann-Whitney U Test(Mann and Whitney, 1947) to demonstrate the statistical significance ($p$-value) of the results.
@@
We implement Locus using the PydanticAI framework (Colvin, 2025). The synthesizer’s tooling is primarily built on top of Multiplier (Goodman, 2025). We leverage the SVF static analysis framework (Sui and Xue, 2016) to perform a lightweight reachability analysis and use KLEE (Cadar et al., 2008) as the symbolic execution engine to perform semantic validation. We use o3-mini-2025-01-31 as the default model with reasoning level set to medium, but we also evaluate other LLMs in Section 4.4.
@@ 4.2. RQ1: Vulnerabilities Reproduction
We apply Locus on both directed fuzzers and coverage-guided fuzzers, comparing the performance with and without Locus. The results are shown in Table 2.
@@ 4.3. RQ2: Cost Analysis
Time cost  In Section 4.2, we focus on measuring TTE, i.e., the time required to trigger the target canary during fuzzing. However, evaluating only the fuzzing phase can be misleading when assessing the overall effectiveness of a directed fuzzer. Many directed fuzzers perform expensive static analysis on the target program before fuzzing begins(Böhme et al., 2017; Huang et al., 2024a, b). Likewise, Locus requires additional preprocessing steps, including codebase indexing, symbolic validation of predicates, and agentic predicate synthesis. Although such preprocessing costs are often treated as a one-time effort amortized over fuzzing, we find that some approaches incur excessive analysis time, sometimes even longer than the time required to discover the bug by the fuzzer.
@@
LLM token cost  We analyze the LLM token usage to assess the financial feasibility of deploying Locus. Table 4 shows the token cost for our Locus workflow. Locus takes 457k tokens (equivalent to \$0.72 USD) to generate predicates for one sample in the Magma benchmark on average. The monetary token costs show the potential for Locus to act as an affordable step in assisting fuzzing.
@@ 4.4. RQ3: Ablations
Table 5. TTEs by ablating different designs and models
@@
Ablate different designs
@@
Ablate different LLMs
@@
with o3-mini 41101 19 1122 682 4961
with Deepseek R1 45104 53 992 1439 5433
with Gemini Flash 2.0 38215 130 1517 2454 5274
@@
with o3-mini 8537 5 4787 7833 6748
with Deepseek R1 23699 6 3562 3518 7023
with Gemini Flash 2.0 7020 12 7206 40332 3921
@@
Table 5 shows that the design choices in Locus consistently improve the performance. Specifically, we observe that while predicates generated under the base setting may occasionally accelerate fuzzing, they are not consistently beneficial. Without validation, Locus sometimes generates false predicates to the program, as evidenced by the cases PNG007, TIF014, and SQL018.